A man is four times as old as his son,in another four years time he will be three times as old as his son,what is their ages if the product of their ages is 42?
Answer by ankor@dixie-net.com(22740)

A man is four times as old as his son, in another four years time he will be three times as old as his son,
what is their ages if the product of their ages is 42?
:
Let m = man's age
Let s = son's age
;
Write an equation for each statement, simplify if possible:
:
"A man is four times as old as his son,"
m = 4s
:
in another four years time he will be three times as old as his son,
m + 4 = 3(s+4)
m + 4 = 3s + 12
m = 3s + 12 - 4
m = 3s + 8
replace m with 4s
4s = 3s + 8
4s - 3s = 8
s = 8 yrs is the son's age
then
4(8) = 32 yrs is the man's age
:
The last statement, just screws up the problem and should be ignored!
: